# Causal reasoning in difference graphs

### v1

@@ Abstract
In epidemiology, understanding causal mechanisms across different populations is essential for designing effective public health interventions. Recently, difference graphs have been introduced as a tool to visually represent causal variations between two distinct populations. While there has been progress in inferring these graphs from data through causal discovery methods, there remains a gap in systematically leveraging their potential to enhance causal reasoning. This paper addresses that gap by establishing conditions for identifying causal changes and effects using difference graphs and observational data. It specifically focuses on identifying total causal changes and total effects in a nonparametric framework, as well as direct causal changes and direct effects in a linear context. In doing so, it provides a novel approach to causal reasoning that holds potential for various public health applications.
@@ 1 Introduction
In epidemiology, it is critical to understand the causal mechanisms that impact health outcomes across diverse populations, which may differ based on geographic location, socioeconomic status, genetic predispositions, or environmental exposures. Recently, difference graphs have been introduced to represent differences between two Structural Causal Models (SCMs) [Pearl, 2009], with each model corresponding to a distinct population. These graphs offer a refined representation to visually compare and analyze the causal variations between these groups. By clearly delineating these causal differences, epidemiologists can tailor public health policies and interventions more effectively to address the unique needs of each population. Moreover, it has been highlighted that such graphs are particularly effective for root cause analysis, as noted in [Malik et al., 2024]. This utility is especially evident within the framework proposed by [Assaad et al., 2023], which identifies root causes by comparing direct effects under normal conditions to those under anomalous conditions.
Recently, there has been a surge of interest in constructing difference graphs directly from data through methods rooted in causal discovery [Wang et al., 2018, Malik et al., 2024, Bystrova et al., 2024]. The simplest approach involves independently discovering the SCM for each of the two populations and then computing the difference graph. However, a more effective strategy has emerged that involves directly constructing the difference graph, bypassing the two SCMs that define it [Wang et al., 2018, Malik et al., 2024, Bystrova et al., 2024]. This approach reduces the need for stringent assumptions commonly associated with causal discovery [Spirtes et al., 2000, Glymour et al., 2019]. Despite these advancements, an evident gap remains in the existing literature: there has been no systematic investigation into how difference graphs themselves can facilitate or enhance the process of causal reasoning. This oversight presents a unique opportunity to explore the potential of difference graphs as a tool not just for representation but also for driving identifiability of causal changes and causal effects.
@@
In the remainder, Section 2 introduces difference graphs along with necessary tools and terminologies. Section 3 details the identifiability results for total effects. Section 4 focuses on the identifiability of direct effects. Section 5 discusses related works and Section 6 concludes the paper.
@@ 2 Difference graphs, causal changes, and causal effects
In the following, uppercase letters represent variables and lowercase letters represent their values, while letters in blackboard bold denote sets.
This study utilizes the framework of structural causal models (SCMs) introduced by [Pearl, 2009]. Formally, an SCM $\mathcal{M}$ is defined as a 4-tuple $(\mathbb{U},\mathbb{V},\mathbb{F},P(\mathbb{U}))$, where $U$ represents a set of exogenous (latent) variables and $V$ denotes a set of endogenous (observed) variables. The set $F$ consists of functions $\{f_{i}\}_{i=1}^{|V|}$, where each endogenous variable is determined by a function $f_{i}\in\mathbb{F}$ that depends on both a subset of exogenous variables and a subset of endogenous variables (which are also called the direct causes or the parents). In a linear SCM, $f_{i}$ is a linear equation where each coefficient is called path coefficient [Wright, 1920, 1921] and it designates a direct effect. The following assumption regarding exogenous variables is made throughout the paper.
@@
Consistent with common practice in causal inference, it is also assumed that each SCM $\mathcal{M}$ induces a causal graph $\mathcal{G}=(\mathbb{V},\mathbb{E})$ which is assumed to be a directed acyclic graph (DAG). For clarity, any causal graph that is a DAG will be referred to as a causal DAG. In this graph, $\mathbb{V}$ represents the set of vertices, each corresponding to a random variable, and $\mathbb{E}$ includes the directed edges, with each edge $X\rightarrow Y$ indicating a causal relationship between $X$ and $Y$. For any graph $\mathcal{G}=(\mathbb{V},\mathbb{E})$, standard terminology is adopted to describe relations to any vertex $Y\in\mathbb{V}$: $Parents(Y,\mathcal{G})$ denotes the set of vertices from which there is a direct edge to $Y$; $Ancestors(Y,\mathcal{G})$ refers to the collective set including $Y$ and the parents of all vertices that can be recursively traced back from $Y$; $Descendants(Y,\mathcal{G})$ comprises all vertices $X$ for which $Y$ is found within $Ancestors(X,\mathcal{G})$. If a path contains the subpath $X\rightarrow W\leftarrow Y$, then $W$ is a collider on the path. A path from $X$ to $Y$ is considered active given a vertex set $\mathbb{W}$, where neither $X$ nor $Y$ are in $\mathbb{W}$, if every non-collider on the path is not in $\mathbb{W}$, and every collider on the path has a descendant in $\mathbb{W}$. Otherwise, the path is blocked by $\mathbb{W}$. It is important to note that many SCMs can induce the same causal DAG, meaning that a causal DAG can be compatible with many probability distributions.
@@
In epidemiology, it is critical to understand the causal mechanisms that impact health outcomes across diverse populations, which may differ based on geographic location, socioeconomic status, genetic predispositions, or environmental exposures. Recently, difference graphs have been introduced as a tool to represent the differences between two SCMs, each corresponding to a distinct population, offering a refined method to visually compare causal variations between these groups.
@@
Note that multiple unordered pairs of SCMs can define the same difference graph. In Definition 1, difference graphs are defined using SCMs rather than causal DAGs because causal variations may not always be structurally evident by comparing the two causal DAGs, since changes are often parametric rather than structural. That said, Figures 1 and 2, feature six examples of difference graphs ((c), (h), (m) in Figure 1 and (c), (f), (k) in Figures 2) along with pairs of causal DAGs on both side of each difference graph that can also define the same difference graph (indicating structural differences between the two populations). This selection was made intentionally in order to make the concept of difference graphs visually clearer and easier to understand. Nonetheless, it is important to remember that the theoretical results of this paper apply even when the difference between the two populations is not structurally apparent. However, all those results are grounded in the following assumption regarding the distributions that are compatible with the difference graph.
@@
As shown in Figure 2, generally, there is no inherent necessity to impose similarity constraints on two causal DAGs and thus difference graphs can contain cycles. However, to make the results of this paper more comprehensible and consistent with recent literature on difference graphs, we initially assume that the two causal DAGs maintain the same topological ordering.
@@
This assumption implies that the difference graph is also a DAG [Wang et al., 2018] and it can hold in many applications, for example, in biological pathways an upstream gene does not generally become a downstream gene in different conditionsn [Wang et al., 2018]. However, in some domain it is concevable to imagine scienarios where causal relations can change direction under different conditions. For example, in epidemiology, obesity often leads to the development of type 2 diabetes in the general population, with no direct feedback from diabetes to obesity. However, in a subpopulation of individuals taking certain antipsychotic medications, the onset of type 2 diabetes can lead to significant weight gain due to the medication’s side effects, without obesity being the original cause. Similarly, in nephrology, diabetes can lead to chronic kidney disease, as high blood sugar damages the kidneys in the general population. However, in patients with end-stage kidney disease, impaired kidney function can cause insulin resistance, leading to diabetes, without diabetes further worsening kidney disease. As a result, for each finding presented in this paper under Assumption 3, we will also provide an alternative result that extends to cases where Assumption 3 is not satisfied.
Recently, there has been increasing interest in discovering difference graphs directly from data, without first discovering the two SCMs that define them [Wang et al., 2018, Malik et al., 2024, Bystrova et al., 2024]. However, a key question that has yet to be explored, but naturally arises, is whether these graphs can provide conditions for estimating causal changes between two populations using observational data. This relates to a technical concept known as identifiability.
Definition 2 (Causal change identifiability using difference graph).
Let $X$ and $Y$ be distinct vertices in a difference graph $\mathcal{D}=(\mathbb{V},\mathbb{E}^{\mid 1-2\mid})$. The causal change $Q$ of $X$ on $Y$ is identifiable in $\mathcal{D}$ if $Q$ is uniquely computable from any unordered pairs observational distributions consistent with $\mathcal{D}$.
@@
This work considers two types of causal changes: total causal change and direct causal change, also referred to as mechanism change [Tian and Pearl, 2001]. The total causal change of $X$ on $Y$ is defined as the difference between the total effect of $X$ on $Y$ in $\mathcal{M}^{1}$ and the total effect of $X$ on $Y$ in $\mathcal{M}^{2}$. In a nonparametric setting, the total effects is a specific functional of the interventional distribution $\Pr(Y=y\mid\text{do}(X=x))$ (where the $do()$ operator represents the intervention) for different values of $x$. However, for simplicity, the total effect is often reffered to as $\Pr(Y=y\mid\text{do}(X=x))$ and by a slight abuse of notation, in the remainder, the total effect will be denoted as $\Pr(y\mid\text{do}(x))$. Similarly, the direct causal change is the difference between the direct effect of $X$ on $Y$ in $\mathcal{M}^{1}$ and the direct effect of $X$ on $Y$ in $\mathcal{M}^{2}$. In a nonparametric setting, the direct effect can be defined as a specific functional of $\Pr(y\mid do(x),do(S_{x,y}))$ for different values of $x$, known as the controlled direct effect [Pearl, 2009], where $S_{x,y}$ represents the set of all observed variables in the system except $X$ and $Y$ [^2] Another approach to defining a direct effect in a nonparametric setting is through the natural direct effect [Pearl, 2001], which relies on counterfactual notions rather than the do() operator. However, since it cannot be defined purely using interventional concepts, it is considered beyond the scope of this paper.. The controlled direct effect of $X$ on $Y$ is not uniquely defined as it can change with respect to the values of $S_{x,y}$ (some variables in $S_{x,y}$ are known as effect modifier). For instance, the direct effect of a pill on thrombosis is likely different for pregnant and non-pregnant women [Pearl, 2009]. However, in a linear SCM, i.e, $\forall f_{i}\in\mathbb{F}$ $f_{i}$, is a linear function, the direct effect, obtained through differentiation, corresponds to the familiar path coefficient $\alpha_{x,y}$ [Wright, 1920, 1921], which, unlike the non-parametric controlled direct effect, is uniquely defined. Therefore, in this work, whenever direct effect is discussed, it will refer specifically to linear SCMs.
@@
Since the total causal change of $X$ on $Y$ is a functional of the total effect, $\Pr(y\mid\text{do}(x))$, and the direct causal change is a functional of the direct effect, $\alpha_{x,y}$, the total causal change is identifiable if the total effect is identifiable, and likewise, the direct causal change is identifiable if the direct effect is identifiable. Therefore, the focus of the remainder of this paper is to establish the identifiability of both the total and direct effects using difference graphs. We begin by providing a general definition of the identifiability of causal effects using difference graphs, which is adapted from the definition given by Pearl [2009] for the identifiability of causal effects using causal DAGs.
@@
A set of variables $\mathbb{W}$ satisfies the back-door criterion relative to the total effect $\Pr(y\mid\text{do}(x))$ in a causal DAG $\mathcal{G}$ if $\forall W\in\mathbb{W}$, $W\not\in Descendants(X,\mathcal{G})$ and $\mathbb{W}$ blocks every path between $X$ and $Y$ which contains an arrow into $X$.
@@
A set of variables $\mathbb{W}$ satisfies the single-door criterion relative to the direct effect $\alpha_{x,y}$ in a causal DAG $\mathcal{G}$ if $\forall W\in\mathbb{W}$, $W\not\in Descendants(Y,\mathcal{G})$ and $\mathbb{W}$ blocks every path between $X$ and $Y$.
Whenever there is a set $\mathbb{W}$ that satisfied the back-door criterion in a causal DAG, $\Pr(y\mid\text{do}(x))$ is given by the adjsutment formula $\sum_{\mathbb{w}}\Pr(y\mid x,\mathbb{w})\Pr(\mathbb{w})$. Whenever there is a set $\mathbb{W}$ that satisfied the single-door criterion in a causal DAG, $\alpha_{x,y}$ is given by the regression coefficient of $X$ on $Y$ after $\mathbb{W}^{anc}$ is partialled out.
@@
Now that all necessary terminologies and notations have been established, the two problems that this work aims to address can be formally presented.
@@
Consider a known difference graph $\mathcal{D}$ defined by two unknown SCMs that respectively induce two unknown causal DAGs. Is $\Pr(y\mid\text{do}(x))$ identifiable in $\mathcal{D}$ by a common back-door?
@@
Consider a known difference graph $\mathcal{D}$ defined by two unknown linear SCMs that respectively induce two unknown causal DAGs. Is $\alpha_{x,y}$ identifiable in $\mathcal{D}$ by a common single-door?

### v2

@@ Abstract
Understanding causal mechanisms across different populations is essential for designing effective public health interventions. Recently, difference graphs have been introduced as a tool to visually represent causal variations between two distinct populations. While there has been progress in inferring these graphs from data through causal discovery methods, there remains a gap in systematically leveraging their potential to enhance causal reasoning. This paper addresses that gap by establishing conditions for identifying causal changes and effects using difference graphs. It specifically focuses on identifying total causal changes and total effects in a nonparametric setting, as well as direct causal changes and direct effects in a linear setting. In doing so, it provides a novel approach to causal reasoning that holds potential for various public health applications.
@@ 1 Introduction
In epidemiology, it is critical to understand the causal mechanisms that impact health outcomes across diverse populations, which may differ based on geographic location, socioeconomic status, genetic predispositions, or environmental exposures. Recently, difference graphs have been introduced to represent differences between two Structural Causal Models (SCMs) [Pearl, 2009], with each model corresponding to a distinct population. These graphs offer a refined representation to visually compare and analyze the causal variations between these groups. By clearly delineating these causal differences, epidemiologists can tailor public health policies and interventions more effectively to address the unique needs of each population. Moreover, it has been highlighted by Malik et al. [2024] that such graphs are also interesting for root cause analysis [Assaad et al., 2023].
Recently, there has been a surge of interest in constructing difference graphs directly from data through methods rooted in causal discovery [Wang et al., 2018, Chen et al., 2023, Malik et al., 2024, Bystrova et al., 2024]. The simplest approach involves independently discovering the SCM for each of the two populations and then computing the difference graph. However, a more effective strategy has emerged that involves directly constructing the difference graph, bypassing the two SCMs that define it. This approach reduces the need for stringent assumptions commonly associated with causal discovery [Spirtes et al., 2000, Glymour et al., 2019]. Despite these advancements, an evident gap remains in the existing literature: there has been no systematic investigation into how difference graphs themselves can facilitate or enhance the process of causal reasoning. This oversight presents a unique opportunity to explore the potential of difference graphs as a tool not just for representation but also for driving identifiability of causal changes and causal effects.
@@
In the remainder, Section 2 introduces difference graphs along with necessary tools and terminologies. Section 3 details the identifiability results for total effects. Section 4 focuses on the identifiability of direct effects. Section 5 presents related works and offers further discussion. Finally, Section 6 concludes the paper. In the following, uppercase letters represent variables and lowercase letters represent their values, while letters in blackboard bold denote sets.
@@ 2 Difference graphs, causal changes, and causal effects
This study utilizes the framework of structural causal models (SCMs) introduced by Pearl [2009]. Formally, an SCM $\mathcal{M}$ is defined as a 4-tuple $(\mathbb{U},\mathbb{V},\mathbb{F},P(\mathbb{U}))$, where $\mathbb{U}$ represents a set of exogenous (latent) variables and $\mathbb{V}$ denotes a set of endogenous (observed) variables. The set $\mathbb{F}$ consists of functions $\{f_{i}\}_{i=1}^{|\mathbb{V}|}$, where each endogenous variable is determined by a function $f_{i}\in\mathbb{F}$ that depends on its corresponding exogenous variable $U_{i}\in\mathbb{U}$ and a subset of endogenous variables (which are also called the direct causes or the parents). In a linear SCM, $f_{i}$ is a linear equation where each coefficient is called path coefficient [Wright, 1920, 1921] and it designates a direct effect. The following assumption regarding exogenous variables is made throughout the paper.
@@
Consistent with common practice in causal inference, it is also assumed that each SCM $\mathcal{M}$ induces a causal graph $\mathcal{G}=(\mathbb{V},\mathbb{E})$ which is assumed to be a directed acyclic graph (DAG). For clarity, any causal graph that is a DAG will be referred to as a causal DAG. In this graph, $\mathbb{V}$ represents the set of vertices, each corresponding to a random variable, and $\mathbb{E}$ includes the directed edges, with each edge $X\rightarrow Y$ indicating a causal relationship between $X$ and $Y$. For any graph $\mathcal{G}=(\mathbb{V},\mathbb{E})$, standard terminology is adopted to describe relations to any vertex $Y\in\mathbb{V}$: $Pa(Y,\mathcal{G})$ denotes the set of vertices from which there is a direct edge to $Y$; $An(Y,\mathcal{G})$ refers to the ancestors of $Y$ which is the collective set including $Y$ and the parents of all vertices that can be recursively traced back from $Y$; $De(Y,\mathcal{G})$ refers to the descendants of $Y$ which comprises all vertices $X$ for which $Y$ is found within $An(X,\mathcal{G})$. If a path contains the subpath $X\rightarrow W\leftarrow Y$, then $W$ is a collider on the path. A path from $X$ to $Y$ is considered active given a vertex set $\mathbb{W}$, where neither $X$ nor $Y$ are in $\mathbb{W}$, if every non-collider on the path is not in $\mathbb{W}$, and every collider on the path has a descendant in $\mathbb{W}$. Otherwise, the path is blocked by $\mathbb{W}$. It is important to note that many SCMs can induce the same causal DAG and that a causal DAG can be compatible with many probability distributions.
@@
Recently, difference graphs have been introduced as a tool to represent the differences between two SCMs, each corresponding to a distinct population, offering a refined method to visually compare causal variations between these groups.
@@
Note that multiple unordered pairs of SCMs can define the same difference graph. In Definition 1, difference graphs are defined using SCMs rather than causal DAGs because causal variations may not always be structurally evident by comparing the two causal DAGs, since changes are often parametric rather than structural. That said, Figures 1 and 2, feature six examples of difference graphs ((c), (h), (m) in Figure 1 and (c), (f), (k) in Figures 2) along with pairs of causal DAGs on both side of each difference graph that can also define the same difference graph (indicating structural differences [^2] The term ”structural differences” refers to ”topological differences between the two causal DAGs”. between the two populations). This selection was made intentionally in order to make the concept of difference graphs visually clearer and easier to understand. Nonetheless, it is important to remember that the theoretical results of this paper apply even when the difference between the two populations is not structurally apparent. However, all those results are grounded in the following assumption regarding the distributions that are compatible with the difference graph.
@@
As shown in Figure 2, generally, there is no inherent necessity to impose similarity constraints on two causal DAGs and thus difference graphs can contain cycles. However, to make the results of this paper more comprehensible and consistent with recent literature on difference graphs, I initially assume that the two causal DAGs have the same topological ordering.
@@
This assumption implies that the difference graph is also a DAG [Wang et al., 2018] and it can hold in many applications, for example, in biological pathways an upstream gene does not generally become a downstream gene in different conditions [Wang et al., 2018]. However, in some domains, it is possible to imagine scienarios where causal relations can change direction under different conditions. For example, in epidemiology, obesity often leads to the development of type 2 diabetes in the general population, with no direct feedback from diabetes to obesity. However, in a subpopulation of individuals taking certain antipsychotic medications, the onset of type 2 diabetes can lead to significant weight gain due to the medication’s side effects, without obesity being the original cause. Similarly, in nephrology, diabetes can lead to chronic kidney disease, as high blood sugar damages the kidneys in the general population. However, in patients with end-stage kidney disease, impaired kidney function can cause insulin resistance, leading to diabetes, without diabetes further worsening kidney disease. As a result, for each finding presented in this paper under Assumption 3, I will also provide an alternative result that extends to cases where Assumption 3 is not satisfied.
@@
Recently, there has been increasing interest in discovering difference graphs directly from data, without first discovering the two SCMs that define them [Wang et al., 2018, Malik et al., 2024, Bystrova et al., 2024]. However, a key question that has yet to be explored, but naturally arises, is whether these graphs can provide conditions for estimating causal changes between two populations using observational data. This work considers two types of causal changes: total causal change and direct causal change, also referred to as mechanism change [Tian and Pearl, 2001]. The total causal change of $X$ on $Y$ is defined as the difference between the total effect of $X$ on $Y$ in $\mathcal{M}^{1}$ and the total effect of $X$ on $Y$ in $\mathcal{M}^{2}$. In a nonparametric setting, the total effects is a specific functional of the interventional distribution $\Pr(Y=y\mid\text{do}(X=x))$ (where the $do()$ operator represents the intervention) for different values of $x$. However, for simplicity, the total effect is often reffered to as $\Pr(Y=y\mid\text{do}(X=x))$ and by a slight abuse of notation, in the remainder, the total effect will be denoted as $\Pr(y\mid\text{do}(x))$. Similarly, the direct causal change is the difference between the direct effect of $X$ on $Y$ in $\mathcal{M}^{1}$ and the direct effect of $X$ on $Y$ in $\mathcal{M}^{2}$. In a nonparametric setting, the direct effect can be defined as a specific functional of $\Pr(y\mid do(x),do(\mathbb{S}_{x,y}))$ for different values of $x$, known as the controlled direct effect [Pearl, 2009], where $\mathbb{S}_{x,y}$ represents the set of all observed variables in the system except $X$ and $Y$ [^3] Another approach to defining a direct effect in a nonparametric setting is through the natural direct effect [Pearl, 2001], which relies on counterfactual notions rather than the do() operator. However, since it cannot be defined purely using interventional concepts, it is considered beyond the scope of this paper.. The controlled direct effect of $X$ on $Y$ is not uniquely defined as it can change with respect to the values of $\mathbb{S}_{x,y}$ (some variables in $\mathbb{S}_{x,y}$ are known as effect modifier). For instance, the direct effect of a pill on thrombosis is likely different for pregnant and non-pregnant women [Pearl, 2009]. However, in a linear SCM, i.e, $\forall f_{i}\in\mathbb{F}$ $f_{i}$, is a linear function, the direct effect, obtained through differentiation, corresponds to the familiar path coefficient $\alpha_{x,y}$ [Wright, 1920, 1921], which, unlike the non-parametric controlled direct effect, is uniquely defined. Therefore, in this work, whenever direct effect is discussed, it will refer specifically to linear SCMs.
@@
Since the total causal change of $X$ on $Y$ is a functional of the total effect, $\Pr(y\mid\text{do}(x))$, and the direct causal change is a functional of the direct effect, $\alpha_{x,y}$, the total causal change is identifiable if the total effect is identifiable, and likewise, the direct causal change is identifiable if the direct effect is identifiable [^4] In some cases, a causal change can be identified even when the corresponding causal effect is not, e.g., if there is no directed path from $X$ to $Y$, the causal change is clearly zero, regardless of whether the causal effect is identifiable.. Therefore, the focus of the remainder of this paper is to establish the identifiability of both the total and direct effects using difference graphs. I begin by providing a general definition of the identifiability of causal effects using difference graphs, which is adapted from the definition given by Pearl [2009] for the identifiability of causal effects using causal DAGs.
@@
A set of variables $\mathbb{W}$ satisfies the back-door criterion relative to the total effect $\Pr(y\mid\text{do}(x))$ in a causal DAG $\mathcal{G}$ if $\forall W\in\mathbb{W}$, $W\not\in De(X,\mathcal{G})$ and $\mathbb{W}$ blocks every path between $X$ and $Y$ which contains an arrow into $X$.
@@
A set of variables $\mathbb{W}$ satisfies the single-door criterion relative to the direct effect $\alpha_{x,y}$ in a causal DAG $\mathcal{G}$ if $\forall W\in\mathbb{W}$, $W\not\in De(Y,\mathcal{G})$ and $\mathbb{W}$ blocks every path between $X$ and $Y$.
Whenever a set $\mathbb{W}$ satisfies the back-door criterion [^5] The back-door criterion is in general not complete. However, when $X$ is a singleton (as considered in this paper), an adjustment set exists iff there is at least one set that satisfies the back-door criterion [Perković et al., 2018]. in a causal DAG, $\Pr(y\mid\text{do}(x))$ is given by the adjsutment formula $\sum_{\mathbb{w}}\Pr(y\mid x,\mathbb{w})\Pr(\mathbb{w})$. Whenever a set $\mathbb{W}$ satisfies the single-door criterion in a causal DAG, $\alpha_{x,y}$ is given by the regression coefficient of $X$ on $Y$ after $\mathbb{W}$ is partialled out.
@@
In addition to examining identifiability by a common back-door (for the total effect) or by a common single-door (for the direct effect), I also consider cases where the total or direct effect is identifiable because it is evident from the graph that the causal effect does not exist. For example, in a DAG where $Y$ causes $X$, the total and direct effects of $X$ on $Y$ are both identifiable: $\Pr(y\mid\text{do}(x))$ reduces to $\Pr(y)$, and $\alpha_{x,y}$ is zero. In such cases, the total or direct effect are referred to as being identifiable trivially. Now that all necessary terminologies and notations have been established, the two problems that this work aims to address can be formally presented.
@@
Consider a difference graph $\mathcal{D}$ defined by two unknown SCMs that respectively induce two unknown causal DAGs. Is $\Pr(y\mid\text{do}(x))$ identifiable in $\mathcal{D}$ trivially or by a common back-door?
@@
Consider a difference graph $\mathcal{D}$ defined by two unknown linear SCMs that respectively induce two unknown causal DAGs. Is $\alpha_{x,y}$ identifiable in $\mathcal{D}$ trivially or by a common single-door?